OpenStudy (hitaro9):

What is the remainder when you divide 4^44 by 19?

OpenStudy (hitaro9):

Most of the remainder problems thus far have been simple (square for difference of 1 etc) but I'm not sure how to tackle this one.

OpenStudy (hitaro9):

Using mod and congruences

OpenStudy (hitaro9):

Is there an easy way to go about this besides guessing and checking numbers?

OpenStudy (kittiwitti1):

There might be, but sadly, I am not very familiar with this process. ^^; Perhaps someone else can assist you? :]

OpenStudy (kittiwitti1):

with this type of equation*

ganeshie8 (ganeshie8):

there are several ways to do this, one easy mehtod that doesn't require much theory is binary exponential algorithm heard of it before ?

OpenStudy (hitaro9):

No I haven't, sorry.

ganeshie8 (ganeshie8):

Notice 44 = 32 + 8 + 4

ganeshie8 (ganeshie8):

4^2 = 16 = -3 `4^4 = (-3)^2 = 9` `4^8 = 9^2 = 5` 4^16 = 5^2 = 6 `4^32 = 6^2 = -2`

ganeshie8 (ganeshie8):

\[4^{44} = 4^{32+8+4} = 4^{32}\cdot 4^8\cdot 4^4 \equiv -2\cdot 5\cdot 9 \equiv 5 \]

OpenStudy (mathmath333):

strange

OpenStudy (hitaro9):

Okay, that's amazing thank you. Can I ask how you reasoned through that though? Like, each step makes sense, but I would have no idea how to replicate that during an exam.

OpenStudy (mathmath333):

i thought fermat little theorem is efficient

ganeshie8 (ganeshie8):

Yep fermat little thm gives the answer fast

ganeshie8 (ganeshie8):

lets do another problem and im sure you will see why this works

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\normalsize \text{one easy way is to apply fermat's little theorem}\hspace{.33em}\\~\\ a^{n-1}\pmod n=1,n\nmid a \hspace{.33em}\\~\\ \dfrac{4^{44} }{19}\hspace{.33em}\\~\\ =\dfrac{4^{19\times 2}\cdot 4^{8} }{19}\hspace{.33em}\\~\\ =\dfrac{1\cdot 4^{8} }{19}\hspace{.33em}\\~\\ =\dfrac{1\cdot 16^{4} }{19}\hspace{.33em}\\~\\ =\dfrac{ (-3)^{4} }{19}\hspace{.33em}\\~\\ =\dfrac{ (9)^{2} }{19}\hspace{.33em}\\~\\ =\dfrac{ 81 }{19}\hspace{.33em}\\~\\ =\dfrac{5}{19} \hspace{.33em}\\~\\ =5 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (hitaro9):

We haven't used Fermat's Little Theorem yet.

OpenStudy (mathmath333):

i m still unable to understand ur computer algorithm

OpenStudy (mathmath333):

and how can it work on every \(\pmod n\) ?

OpenStudy (mathmath333):

note that \(n\) id prime is my answer

OpenStudy (hitaro9):

And ganeshie, that was the last problem of its kind. There's one last proof dealing with mods that I'm not sure how to tackle if ya want to be an incredible human being again and explain

OpenStudy (hitaro9):

It's different, but also pretty complicated. I need to prove that 4^n is congruent to 1+3n (mod 9)

OpenStudy (hitaro9):

I feel like with a lot of these mod problems I'm able to look at the answer once it's given and be like "oh okay, every step makes sense. No idea how they got from start to finish without blind guessing though"

OpenStudy (hitaro9):

Thank you, bookmarked it for later.

OpenStudy (hitaro9):

I don't have a ton of time left before I've got to drive to classes though, and I still feel like I have no idea how to proceed with this proof. Do you have any ideas?

OpenStudy (mathmath333):

for fermat liite theorm ?

OpenStudy (hitaro9):

You and Gane have been amazing by the way, sorry to have needed so much help.

ganeshie8 (ganeshie8):

It should be easy using binomial thm. Just write 4 as 3+1 and expand

OpenStudy (hitaro9):

No uhh "If Prove 4n ≡ 1 + 3n (mod 9) for any integer n ≥ 1"

OpenStudy (hitaro9):

Prove 4n is congruent to 1+ 3n mod 9

OpenStudy (hitaro9):

Okay 1 second gane

OpenStudy (mathmath333):

what is n here ?

OpenStudy (hitaro9):

A number greater than 1

OpenStudy (hitaro9):

Integer*

ganeshie8 (ganeshie8):

\[\begin{align}4^n = (1+3)^n &= \binom{n}{0} 1^n3^0 + \binom{n}{1} 1^{n-1}3^1 + 3^2(\text{stuff}) \\~\\ &= 1 + 3n + 9(\text{stuff}) \\~\\ &\equiv 1+3n + 0 \pmod{9} \end{align}\]

OpenStudy (mathmath333):

oh its exponent

OpenStudy (hitaro9):

Oh sorry. The copy paste kinda failed the second time. Yeah. It's 4^n is congruent to 3n+1 mod 9

OpenStudy (mathmath333):

perfect!

OpenStudy (hitaro9):

I haven't seen that (n 0 ) Notation. But if I understand the proof I'm going to have to show that 3^n = 3n mod 9 and it basically cylces for values of n?

OpenStudy (hitaro9):

(not at all very well written, but does that sound close to right?)

ganeshie8 (ganeshie8):

familiar with binomial thm ?

OpenStudy (hitaro9):

No, not really.

ganeshie8 (ganeshie8):

Binomial theorem is the first thing to learn in number theory http://en.wikipedia.org/wiki/Binomial_theorem

ganeshie8 (ganeshie8):

It is so fundamental and important.. Most proofs and concepts in number theory depend on / use binomial thm

OpenStudy (hitaro9):

I checked the references, and binomial theorem isn't show in our text until a couple chapters down the line, and our professor hasn't discussed it yet, so for now I have to make due without it.

ganeshie8 (ganeshie8):

makes no sense to study congruences w/o covering binomial thm first

OpenStudy (mathmath333):

yes binomial is very important

ganeshie8 (ganeshie8):

(im just saying that to stress the importance of binomila thm haha!)

OpenStudy (mathmath333):

i said beacause that is really cool and important as i never studied mod arithmatic , just used applications of binomial through out

ganeshie8 (ganeshie8):

Indeed it is cool and wont take much time to learn/use it

ganeshie8 (ganeshie8):

But you will be amazed by the complicated problems that it allows you to solve.. !

OpenStudy (hitaro9):

Alright, I'll be sure to look in to it later tonight.

ganeshie8 (ganeshie8):

What textbook are you using ?

OpenStudy (hitaro9):

Elementary Number Theory and its Applications - Kenneth H Rosen 6th edition

OpenStudy (hitaro9):

Just to make sure, I basically need to show 3^n = 3n mod 9 right?

ganeshie8 (ganeshie8):

thats not true, try n=2

ganeshie8 (ganeshie8):

3^2 = 3*2 mod 9 9 = 6 mod 9 0 = 6 mod 9 false

OpenStudy (hitaro9):

Okay right. Sorry

ganeshie8 (ganeshie8):

you need to know binomial thm to understand above proof... let me see if we can avoid it :)

ganeshie8 (ganeshie8):

we may try induction

OpenStudy (hitaro9):

Induction seems odd. Or maybe I just don't understand cause I'm new at this, but doesn't it change for every n+1?

ganeshie8 (ganeshie8):

i didnt get that.. what do you mean by change for every n+1 ?

OpenStudy (hitaro9):

Umm. Nevermind. You go ahead and explain and I'll see if I don't follow anything.

ganeshie8 (ganeshie8):

Induction is pretty straightforward and simple as we always know where to start

OpenStudy (kittiwitti1):

I'm only here 'cos notifications will break my computer. ; o;

OpenStudy (hitaro9):

It's cool Kitti. I got to go in 15 minutes so hopefully this isn't on too much longer.

OpenStudy (kittiwitti1):

Alright xD

ganeshie8 (ganeshie8):

To prove : \(4^n \equiv 3n+1 \pmod{9}\) Base case : \(4^1 = 4 = 3*1+1 \pmod{9}~\color{green}{\checkmark}\) Induction step : Assume \(4^k\equiv 3k+1\pmod{9}\) is true. \[4^{k+1}\equiv 4\cdot 4^k\equiv 4(3k+1) = 12k+4 \equiv 3k+4 = 3(k+1)+1\pmod{9} \] \(\blacksquare\)

OpenStudy (hitaro9):

Okay, that's incredible. You're amazing, thank you so much.

OpenStudy (kittiwitti1):

ganeshie = the Math God of OS

OpenStudy (kittiwitti1):

and hartnn ofc

ganeshie8 (ganeshie8):

\(\large \color{magenta}{\heartsuit }\)

OpenStudy (kittiwitti1):

:)

OpenStudy (hitaro9):

Okay, again, thank you so much Gane and mathmath for all your help today. You both are incredible